### v1

@@ 1 Introduction
Figure 1: (a, b, c) Our proposed CityGS achieves the SOTA rendering fidelity on Small City scene (5620 training images, 740 test images) of MatrixCity dataset. The setting of baseline 3DGS is discussed in Sec. 4.1. (d, e) Here Z denotes camera height. Without LoD, CityGS would render over 20 million points, leading to considerable VRAM and time costs. The LoD saves VRAM and enables real-time performance under various scales. Note that the FPS is tested with CUDA synchronization for objective evaluation.
@@ 3.2.2 Primitives and Data Division
Considering the irregular distribution present in most scenes, direct block partitioning often results in uneven Gaussian primitives assignment. For instance, in the leftmost section of Fig. 2, the left bottom grid is assigned with a few Gaussians, while the central grid is densely occupied. As a result, under the same training schedule, the sparse blocks tend to overfit while the dense blocks may lack sufficient training. To address this issue, we first contract the Gaussians to a bounded cubic region. Assume the predefined foreground region, i.e. the pink square in Fig. 2, is depicted by the minimum and maximum positions of its corners $\boldsymbol{p}_{\min}$ and $\boldsymbol{p}_{\max}$. We normalize Gaussian positions as $\hat{\boldsymbol{p}}_{\boldsymbol{k}}=2\left(\boldsymbol{p}_{\boldsymbol{k}}-\boldsymbol{p}_{\min}\right)/\left(\boldsymbol{p}_{\max}-\boldsymbol{p}_{\min}\right)-1$. Consequently, the positions of foreground Gaussians fall within the range $[-1,1]$. The subsequent contraction is executed using the following function [37]:
@@ 4.1.2 Implementations and Baselines
Our method first trains coarse global Gaussian priors for further block-wise refinement. The training of priors for Sci-Art, Residence and Rubble adheres to the parameter settings outlined in 3DGS [9]. But for Building and MatrixCity, we halve the learning rate of position and scaling to prevent underfitting caused by aggressive optimization. In the fine-tuning phase, we train each block for another 30,000 iterations with coarse global Gaussian prior as initialization. Furthermore, the learning rate of position is reduced by 60%, while that of scaling is reduced by 20%, compared to setting in 3DGS [9]. The detailed foreground range and block dimensions can be found in the Appendix. Our method is benchmarked against Mega-NeRF [33], Switch-NeRF [46], GP-NeRF [45], and 3DGS [9]. Since the datasets contain thousands of images, which are much larger than that used in the original 3DGS paper [9], we adjust the total iteration to 60,000, while densifying Gaussians from iteration 1,000 to 30,000 with an iteration interval of 200. This strong baseline is named as 3DGS. In LoD, we evaluate on the MatrixCity dataset. We use 3 detail levels, where LoD 2 is the finest and LoD 0 is the coarsest. $n_{MAD}$ mentioned in Sec. 3.3.2 is set to 4. The blocks within 0m to 200m are represented by LoD 2, blocks within 200m to 400m are represented by LoD 1, while others are represented by LoD 0.
@@ 4.2 Comparison with SOTA
For an apple-to-apple rendering quality comparison with the SOTA large-scale scene reconstruction strategy, we use the performance of the CityGS with no LoD. The quantitative result is shown in Table 1. Due to page limits, we put results of Sci-Art in the Appendix. It can be observed that our method outperforms the NeRF-based baselines by a large margin. On the one hand, as far as we know, our method is the first of all attempts that successfully reconstructs the whole MatrixCity with highly variable camera altitude, ranging from 150m to 500m [13]. The PSNR arrives at 27.46 and qualitative results presented in Fig. 5 also validate the high fidelity of our renderings. Compared with the baseline 3DGS, our CityGS can capture much richer details. More visualizations on MatrixCity can be found in the Appendix. On the other three realistic scenes, our method enables much higher SSIM and LPIPS, which indicates outstanding visual quality. As shown in Fig. 4, thin structures like girder steel and window frames can be well reconstructed. The details of complex structures like grass and rubble are well recovered as well.
@@ A Additional Experimental Results
Besides, as shown in Table S1, LoD significantly improves efficiency, especially for extremely large-scale scenes such as MatrixCity. Compared with our CityGS, the 3DGS [9] possesses faster speed but significantly lower rendering quality. The main reason is that the original 3DGS requires sufficiently large iterations and memory to optimize the whole scene with thousands of images. Bounded by computation resources, the capacity of the trained original 3DGS is too limited to well represent the whole large-scale scene.

### v2

@@ 1 Introduction
Figure 1: (a, b, c) Our proposed CityGS achieves the SOTA rendering fidelity on Small City scene (5620 training images, 740 test images) of MatrixCity dataset. The setting of baseline ${}^{\dagger}$ is discussed in Sec. 4.1. (d, e) Here Z denotes camera height. Without LoD, CityGS would render over 20 million points, leading to considerable VRAM and time costs. The LoD saves VRAM and enables real-time performance under various scales. Note that the FPS is tested with CUDA synchronization for objective evaluation.
@@ 3.2.2 Primitives and Data Division
Under normally adopted training view distribution, the central part of the considered region receives more supervision and is thus represented by dense Gaussians. In contrast, the boundary regions are sparsely supervised, and thus are occupied by sparsely and randomly distributed Gaussians. A direct Gaussian partitioning according to a uniform grid will lead to considerably uneven Gaussian assignment. For instance, in the leftmost section of Fig. 2, the left bottom grid is assigned with a few Gaussians, while the central grid is densely occupied. As a result, under the same training schedule, the boundary blocks prone to overfit the sparse supervision, while the central blocks may underfit the dense supervision.
To address this issue, we first contract the Gaussians to a bounded cubic region. The contraction defines a linear mapping region. This region is named the foreground region, i.e. the pink square in Fig. 2. It is depicted by the minimum and maximum positions of its corners $\boldsymbol{p}_{\min}$ and $\boldsymbol{p}_{\max}$. The Gaussians outside this region will experience nonlinear coordinate transformation. Specifically, We first normalize Gaussian positions as $\hat{\boldsymbol{p}}_{\boldsymbol{k}}=2\left(\boldsymbol{p}_{\boldsymbol{k}}-\boldsymbol{p}_{\min}\right)/\left(\boldsymbol{p}_{\max}-\boldsymbol{p}_{\min}\right)-1$. Consequently, the positions of foreground Gaussians fall within the range $[-1,1]$. The subsequent contraction is executed using the following function [37]:
@@ 4.1.2 Implementations and Baselines
Our method first trains coarse global Gaussian priors for further block-wise refinement. The training of priors for Sci-Art, Residence and Rubble adheres to the parameter settings outlined in 3DGS [9]. But for Building and MatrixCity, we halve the learning rate of position and scaling to prevent underfitting caused by aggressive optimization. In the fine-tuning phase, we train each block for another 30,000 iterations with coarse global Gaussian prior as initialization. Furthermore, the learning rate of position is reduced by 60%, while that of scaling is reduced by 20%, compared to setting in 3DGS [9]. The detailed foreground range and block dimensions can be found in the Appendix. Our method is benchmarked against Mega-NeRF [33], Switch-NeRF [46], GP-NeRF [45], and 3DGS [9]. Since the datasets contain thousands of images, which are much larger than that used in the original 3DGS paper [9], we adjust the total iteration to 60,000, while densifying Gaussians from iteration 1,000 to 30,000 with an iteration interval of 200. This strong baseline is named as ${}^{\dagger}$. In LoD, we evaluate on the MatrixCity dataset. We use 3 detail levels, where LoD 2 is the finest and LoD 0 is the coarsest. $n_{MAD}$ mentioned in Sec. 3.3.2 is set to 4. The blocks within 0m to 200m are represented by LoD 2, blocks within 200m to 400m are represented by LoD 1, while others are represented by LoD 0.
@@ 4.2 Comparison with SOTA
For an apple-to-apple rendering quality comparison with the SOTA large-scale scene reconstruction strategy, we use the performance of the CityGS with no LoD. The quantitative result is shown in Table 1. Due to page limits, we put results of Sci-Art in the Appendix. It can be observed that our method outperforms the NeRF-based baselines by a large margin. On the one hand, as far as we know, our method is the first of all attempts that successfully reconstructs the whole MatrixCity with highly variable camera altitude, ranging from 150m to 500m [13]. The PSNR arrives at 27.46 and qualitative results presented in Fig. 5 also validate the high fidelity of our renderings. Compared with the baseline ${}^{\dagger}$, our CityGS can capture much richer details. More visualizations on MatrixCity can be found in the Appendix. On the other three realistic scenes, our method enables much higher SSIM and LPIPS, which indicates outstanding visual quality. As shown in Fig. 4, thin structures like girder steel and window frames can be well reconstructed. The details of complex structures like grass and rubble are well recovered as well.
@@ A Additional Experimental Results
Besides, as shown in Table S1, LoD significantly improves efficiency, especially for extremely large-scale scenes such as MatrixCity. Compared with our CityGS, the ${}^{\dagger}$ [9] possesses faster speed but significantly lower rendering quality. The main reason is that the original 3DGS requires sufficiently large iterations and memory to optimize the whole scene with thousands of images. Bounded by computation resources, the capacity of the trained original 3DGS is too limited to well represent the whole large-scale scene.